• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E. Burtin

    Log entry time 17:00:44 on October15,2005

    Entry number 156954

    keyword=Access + trip of the R septum

    Ed is looking at what could be causing the problems on the right arm spectrometer.
    David is investigating some missing signal on the compton polarimeter
    rebooted the magnet control.
    Bryan fixed the "sync" in the right arm.
    trip of the septum and of most of the magnets.